# The Dynamics of Collective Creativity in Human-AI Social Networks

### v1

@@ Abstract
Generative AI is reshaping modern culture, enabling individuals to create high-quality outputs across domains such as images, text, and music. However, we know little about the impact of generative AI on collective creativity. This study investigates how human-AI interactions shape collective creativity within experimental social networks. We conducted large-scale online experiments with 879 participants and AI agents in a creative writing task. Participants (either humans or AI) joined 5×5 grid-based networks, and were asked to iteratively select, modify, and share stories. Initially, AI-only networks showed greater creativity (rated by a separate group of 94 human raters) and diversity than human-only and human-AI networks. However, over time, hybrid human-AI networks became more diverse in their creations than AI-only networks. In part, this is because AI agents retained little from the original stories, while human-only networks preserved continuity. These findings highlight the value of experimental social networks in understanding human-AI hybrid societies.
@@ Introduction
From cave paintings to symphonies, creativity has defined the human experience – our ability to imagine new possibilities and bring them to life. Today, we stand at a fascinating turning point as generative artificial intelligence (AI) – or AI agents – are transforming the creative process by which humans formulate ideas and put them into practice (?, ?). These AI agents are not merely assisting tools but active partners in the creative process (?, ?, ?), collaborating with writers, musicians, and artists in unprecedented ways. In this hybrid society, networks of interacting humans and intelligent machines constitute complex social systems for which the quality of the collective outcomes cannot be deduced from either human or AI behaviour alone (?, ?). However, studying how collective creativity emerges in hybrid networks of humans and AI agents remains a key challenge.
While generative AI has been shown to increase the quality and efficiency of routine tasks – such as customer support, programming, and academic writing (?, ?), its impact on human creativity remains poorly understood. Some evidence suggests that human-AI collaboration may enhance individual’s creativity (?, ?, ?), increasing the efficiency and speed of generating new ideas. In contrast, other evidence suggests that aesthetic and cultural biases embedded in generative AI models can limit global diversity, leading to homogenization effects in art and culture (?, ?, ?). Although this research has begun to demonstrate the potential impacts of generative AI on human creativity, it has thus far been difficult to quantify how much of human creative potential is driven by specific interactions, and how these can be enhanced by human-AI collaboration. In part, this is because studying these processes in controlled and scalable experimental settings is challenging.
Here, we developed an automated experimental pipeline to study collective creativity in social networks involving humans and generative AI, building on our recent method for integrating human participants within simulated experimental social networks (?, ?). In each “generation” of the experiment, participants (either human or AI) select and modify creative stories from their network neighbours and share their creations with participants in the next generation (Figure 1A). Over time, features of the collective creations emerge and evolve (e.g., quality, diversity), allowing us to establish causal links between experimental conditions – such as different levels of human-AI interaction – and the emergence of collective creativity. Specifically, we systematically compare (1) human-only, (2) AI-only, and (3) human-AI collaboration within the same simulated networks (Figure 1B). We focus on storytelling, an iterative cognitive process (?, ?, ?) central to many creative problems in the real-word, such as scientific writing, entertainment, marketing, and design. This creative modality has substantial opportunities for generative AI to collaboratively ideate and create with humans, for example, by helping to brainstorm new ideas, generate storylines, and improve their writing style and tone (?, ?, ?).
Our approach offers valuable insights into the conditions that enable effective human-AI collaboration and the dynamics of these interactions — particularly how diversity and creativity emerge and shift over iterative creative processes. More broadly, this work highlights the potential of real-world experiments with ecologically valid, open-ended tasks as a powerful approach for studying human-AI collective intelligence and creativity.
@@
Yet, how effective are AI tools? Despite increasing investment in development, research on AI-assisted decision-making presents mixed findings. While some studies indicate that AI enhances performance in domains such as healthcare, customer service, and scientific research (?, ?, ?), others suggest the opposite (?, ?, ?). A recent systematic review and meta-analysis of 106 experimental studies (Vaccaro et al., 2024) confirmed this divide, though, on average, human-AI teams performed better than humans working alone. One promising approach to improving human-AI collaboration is to conceptualize AI as a “thought partner”—an active participant in cognitive processes that engages in reciprocal dialogue with human users (Collins et al., 2024). Unlike traditional AI tools, machine “thought partners” dynamically infer human intentions and internal states, adapting their responses in ways that mirror human-to-human interactions, thereby fostering deeper and more productive collaboration.
Human-AI Collaboration
The integration of AI into human society has far-reaching implications, particularly at the collective level. AI influences not just individuals but also the broader dynamics of social systems (?, ?). Thus, it is becoming incredibly important to understand the role of generative AI in shaping collective behavior, focusing on the emergent structures that arise from interactions between multiple individuals and AI agents simultaneously (?, ?, ?).
@@
Human creativity research has extensively used psychometric measures of creative processes for decades (?, ?). Traditional “creativity tests” – such as the Structure of the Intellect divergent production tests (?, ?) and Tests of Creative Thinking (?, ?) – remain in wide use in creativity research to this day. These tests rely on simple measures of association, such as generating lists of non-repeating words or brainstorming novel uses for a given object. Advances in computational models of semantic networks provide powerful solutions to address limitations in previous research, allowing researchers to quantitatively examine how people navigate a semantic space of concepts and revealing key search strategies associated with creativity (?, ?). Moreover, some studies have explored open-ended creative tasks that do not rely on language, such as visual art creation on grid-based images, modeling the balance between exploration and exploitation in creative processes using computational approaches (?, ?, ?, ?). However, this research primarily treats creativity at the individual level and does not consider AI as an integral part of the creative process.
Collective Creativity in Human-AI Hybrid Societies
The introduction of AI is fundamentally reshaping human creativity (?, ?), where AI agents not only collaborate with individuals but also shape the social environments in which humans operate (?, ?). This nonlinear influence presents a significant challenge for studying human-AI interactions at a collective level. ? (?) demonstrated that while generative AI can enhance individual creativity, it may simultaneously reduce collective diversity, the range and originality of ideas produced by a group. Although this study provides an important initial look at the collective consequences of AI-driven creativity, it did not capture how these effects evolve among diverse agents interacting over time within social networks.
Our approach addresses these limitations by introducing a method that is both ecologically valid and open-ended, enabling the study of large-scale social interactions involving hundreds of participants within social networks. Here, we leveraged our recent capability to design experiments that incorporate real human participants within experimental social networks (?, ?). By using iterative storytelling as a creative task, we examine how ideas propagate, transform, and diversify in networks the involve humans, AI, and hybrid human-AI collaborations. This framework allows us to analyze the complex interplay between human and AI agents in dynamic creative processes at a collective level.
@@
We conducted large-scale online experiments on collective creativity, where participants (either human or AI) were embedded in a 5$\times$5 social network and engaged in a creative storytelling task. Participants selected and modified stories from their neighbours in the netowrk, and their creations were presented to other participants over 25 iterations (Figure 1A). The same prompt was used for both human and AI agents: Please creatively elaborate on the story, adding your own details and ideas”. Participants were randomly assigned to one location (node) in the network and only contributed once to the experiment. All networks were initialized with the same story:
@@
We compared three experimental conditions (Figure 1B): (1) human-only, where all nodes were occupied by human participants (N=625), (2) AI-only, where all nodes were simulated using GPT-4o (625 calls to the OpenAI API, and (3) human-AI, consisting of an equal distribution of human participants (N=254, 50%) and GPT-4o agents (50%).
@@
To asses the quality of creations across conditions, we conducted a validation study with a separate group of 100 human participants. Each participant was presented with a randomized selection of stories from all experimental conditions and asked to rate their creativity on a 5-point scale, ranging from 1 (not creative at all) to 5 (extremely creative). Each participant evaluated 20 different stories. Crucially, participants were not informed whether stories were created by humans or AI.
In addition to subjective creativity, we measured the diversity of stories by computing their semantic similarity. Specifically, we embedded each story using a TF-IDF vectorization approach and calculated the pairwise cosine similarity between all stories within predefined iteration groups. Diversity was operationalized as the inverse of the mean cosine similarity, with lower similarity indicating greater diversity.
@@ Results
To visualize the difference between human and AI-generated stories, we created a semantic embedding space based on all written stories. Figure 3 shows the UMAP embedding of the stories, visualizing the dynamics between the three conditions. Each point represents a single story, and the proximity of points reflects their semantic similarity in the original high-dimensional space. To compute the high-dimensional space, we first generated semantic embeddings for each story using a pre-trained model (all-MiniLM-L6-v2; ?, ?). We then applied UMAP to reduce these embeddings to two dimensions, allowing us to visualize the semantic similarities among the stories.
@@ Discussion
A crucial consideration in interpreting our results is the nature of the creative task itself. The storytelling task used in this study aligns with those traditionally employed in human creativity research (?, ?, ?). However, the task proved to be relatively easy for a large language models like GPT, which consistently outperformed humans alone both in creativity and diversity. This does not necessarily imply that AI is intrinsically more creative than humans; rather, it highlights the limitations of using tasks designed for human cognition as benchmarks for AI creativity. AI’s advantage may stem from its ability to rapidly generate highly novel yet semantically coherent text, a skill optimized by its training on vast datasets (?, ?, ?). Conversely, human creativity and creative writing is influenced by memory constraints, cognitive biases, and adherence to implicit narrative conventions (?, ?).
@@
Our study aimed to simulate social-naturalistic dynamics within experimental societies. However, our approach diverges from how ecologically valid social interactions unfold in the real world. Unlike our experiments, where participants engaged only once, individuals in real-world settings continuously interact with each other over time. Additionally, real-world exchanges are far more complex than the constrained task of creating simple stories used in our experiment. Another key limitation is the static structure of our simulated social network, which used a grid-based topology. Real social networks, such as groups of friends or work relationships, tend to have a modular structure with irregular patterns and evolving connections, allowing individuals to actively shape their interactions by forming and adjusting their social ties rather than being confined to static connections. Finally, our paradigm avoids any direct communication between participants and masks agent information, such as whether your partner is a human or an AI.

### v3

# The Dynamics of Collective Creativity in Human-AI Hybrid Societies

@@ Abstract
Generative AI is shaping an increasingly hybrid society, where ideas and cultural artefacs are created both by humans and intelligent machines. Human creativity is influenced in complex, nonlinear ways by the actions of AI-driven agents within their social networks, but these influences are difficult to measure using traditional methods. This study examines how human-AI interactions shape the evolution of collective creation within large-scale social network experiments, where human and AI participants collectively create stories. Participants (either humans or AI) joined 5×5 grid-based networks in which stories were selected, modified, and shared over many iterations. Initially, AI-only networks showed greater creativity (rated by a separate group of human raters) and collective diversity of stories than human-only and human-AI networks. However, over time, hybrid human-AI networks became more diverse in their creations than AI-only networks. In part, this is because AI agents retained little from the original stories, while human-only networks preserved continuity. These findings highlight the value of experimental social networks in understanding human-AI hybrid societies.
@@ Introduction
From cave paintings to symphonies, creativity has defined the human experience – our ability to imagine new possibilities and bring them to life. Today, we stand at a fascinating turning point as generative artificial intelligence (AI) is transforming the creative process by which humans formulate ideas and put them into practice (?, ?). AI models are not merely assisting tools but active partners in the creative process (?, ?, ?), collaborating with writers, musicians, and artists in unprecedented ways. In this hybrid society, networks of interacting humans and intelligent machines constitute complex social systems for which the quality of the collective outcomes cannot be deduced from either human or AI behaviour alone (?, ?). However, studying how collective creativity emerges in hybrid networks of humans and AI agents remains a key challenge.
While generative AI has been shown to increase the quality and efficiency of routine tasks – such as customer support, programming, and academic writing (?, ?), its impact on human creativity remains poorly understood. Some evidence suggests that human-AI collaboration may enhance individual’s creativity (?, ?, ?), increasing the efficiency and speed of generating new ideas. In contrast, at the same time aesthetic and cultural biases embedded in generative AI models can limit global diversity, leading to homogenization effects in art and culture (?, ?, ?). This research demonstrates the potential impacts of generative AI on human creativity, but we still know little about how good ideas emerge and evolve within human-AI social networks, and the mechanisms that enable effective collaboration between humans and machines.
Our approach introduces a method that is both ecologically valid and open-ended, enabling the study of large-scale social interactions involving hundreds of participants within social networks. Here, we leveraged our recent capability to design experiments that incorporate real human participants within experimental social networks (?, ?). By using iterative storytelling as a creative task, we examine how ideas propagate, transform, and diversify in networks the involve humans, AI, and hybrid human-AI collaborations. This framework allows us to analyze the complex interplay between human and AI agents in dynamic creative processes at a collective level.
@@
Yet, how effective are AI tools? Despite increasing investment in development, research on AI-assisted decision-making presents mixed findings. While some studies indicate that AI enhances performance in domains such as healthcare, customer service, and scientific research (?, ?, ?), others suggest the opposite (?, ?, ?). A recent systematic review and meta-analysis of 106 experimental studies (Vaccaro et al., 2024) confirmed this divide, though, on average, human-AI teams performed better than humans working alone.
Rather than viewing AI systems merely as assisting tools, a more promising conceptual shift—drawn from the science of collaborative cognition—is to understand them as ”partners in thought”: systems designed to be reasonable, insightful, knowledgeable, and trustworthy, capable of actively thinking and creating with humans (?, ?). This perspective moves beyond the instrumental view of AI as passive aids and toward recognizing their potential as cognitive collaborators. As AI models continue to advance and proliferate, they will become increasingly integrated into human interactions and everyday activities, participating in complex, multimodal problem-solving and contributing to collaborative creative processes.
AI and Collective Dynamics
As AI systems become embedded in everyday human activity as collaborative partners, their impact extends beyond individuals to influence collective behavior and social systems at large (?, ?, ?, ?). Thus, it is becoming incredibly important to understand the role of generative AI in shaping collective behavior, focusing on the emergent structures that arise from interactions between multiple individuals and AI agents simultaneously (?, ?).
@@
AI and Collective Creativity
Creativity is typically assessed along two core dimensions: novelty and usefulness (?, ?, ?). Novelty captures the extent to which an idea deviates from existing conventions or expectations, while usefulness reflects the idea’s practicality, coherence, or relevance, such as whether a short story could plausibly be published as a book. Beyond these individual-level criteria, creativity can also be understood at the collective level, where it reflects the diversity of ideas produced across a group or system (?, ?, ?, ?). In this view, a creative ecosystem is not just one that generates isolated novel ideas, but one that fosters a wide range of distinct and meaningful outputs.
Generative AI can impact creativity in at least two ways. AI can enhance creativity by acting as a cognitive catalyst, suggesting unexpected ideas, combining concepts in novel ways, and expanding the exploratory space available to human creators (?, ?, ?). Conversely, AI can hinder creativity by promoting convergence and homogenization (?, ?, ?). Understanding when and how AI enhances versus constrains collective diversity is thus critical for assessing its long-term impact on human cultural production. ? (?) demonstrated that while generative AI can enhance individual creativity, it may simultaneously reduce collective diversity, the range and originality of ideas produced by a group. This study provides an important initial look at the collective consequences of AI-driven creativity, but it did not capture how these effects evolve among diverse agents interacting over time within social networks.
@@
Creativity has been studied extensively in psychology and computer science. In psychology, human creativity has been investigated using psychometric measures developed over decades (?, ?). Classic “creativity tests”—such as the Structure of the Intellect divergent production tasks (?, ?) and the Torrance Tests of Creative Thinking (?, ?)—remain widely used today. These assessments typically involve language-based tasks, such as listing unusual uses for a common object or generating non-repeating words. More recent advances leverage computational models of semantic networks to address the limitations of these tests. These models allow researchers to quantitatively trace how individuals navigate conceptual spaces, revealing cognitive search strategies associated with creative thought (?, ?). Beyond verbal tasks, some studies have explored non-linguistic forms of creativity—such as visual art production in grid-based environments—to model how people balance exploration and exploitation in open-ended tasks using computational techniques (?, ?, ?, ?). However, this research primarily treats creativity at the individual level and does not consider AI as an integral part of the creative process
In computer science, creativity has traditionally been approached from the perspective of computational creativity, a field that explores how machines can be programmed to perform tasks typically associated with human creative behavior (?, ?, ?). This includes the development of algorithms that can generate music, visual art, and poetry. Methods range from rule-based systems and evolutionary algorithms to neural networks and large-scale generative models. A central aim is to build systems that exhibit properties of creativity—such as novelty, value, and surprise (?, ?, ?). Recent advances in generative AI, particularly large language and image models, have expanded the scope of machine-generated content, making it possible to simulate high-level creative tasks with increasing fluency and realism (?, ?, ?). However, this work often emphasizes output quality rather than the cognitive and social processes underlying human-AI creativity, and tends to focus on individual agents rather than collective dynamics. Integrating insights from psychology with computational methods offers a promising path forward for understanding creativity as a collective process that emerges in networks of interacting humans and machines.
@@
We conducted large-scale online experiments on collective creativity, where participants (either human or AI) were embedded in a 5$\times$5 social network and engaged in a creative storytelling task. Participants selected and modified stories from their neighbours in the network, and their creations were presented to other participants over 25 iterations (Figure 1A). The same prompt was used for both human and AI agents: Please creatively elaborate on the story, adding your own details and ideas”. Participants were randomly assigned to one location (node) in the network and only contributed once to the experiment. All networks were initialized with the same story:
@@
We compared three experimental conditions (Figure  1B): (1) human-only, where all nodes were occupied by human participants (N=625), (2) AI-only, where all nodes were simulated using GPT-4o (625 calls to the OpenAI API, and (3) human-AI, consisting of an equal distribution of human participants (N=254, 50%) and GPT-4o agents (50%). To minimize bias, participants were not informed that any of the stories may have been generated by AI agents.
@@
To assess the quality of creations across conditions and iterations, we conducted a validation study with a separate group of 100 human participants. Each participant was presented with a randomized selection of stories from all experimental conditions and iterations and asked to rate their creativity on a 5-point scale, ranging from 1 (not creative at all) to 5 (extremely creative). Each participant evaluated 20 different stories. Crucially, participants were not informed whether stories were created by humans or AI.
In addition to subjective creativity, we measured the diversity of stories by computing their semantic similarity. Specifically for the diversity analysis, we embedded each story using a TF-IDF vectorization approach (?, ?) and calculated the pairwise cosine similarity between all stories within predefined iteration groups. Diversity was operationalized as the inverse of the mean cosine similarity, with lower similarity indicating greater diversity.
@@ Results
To visualize the difference between human and AI-generated stories, we created a semantic embedding space based on all written stories. Figure 3 shows the UMAP embedding of the stories, visualizing the dynamics between the three conditions. Each point represents a single story, and the proximity of points reflects their semantic similarity in the original high-dimensional space. To compute the high-dimensional space used for the UMAP projection, we first generated semantic embeddings for each story using a pre-trained transformer model (all-MiniLM-L6-v2; ?, ?). We then applied UMAP to reduce these embeddings to two dimensions, allowing us to visualize the semantic similarities among the stories.
@@ Discussion
A crucial consideration in interpreting our results is the nature of the creative task itself. The storytelling task used in this study aligns with those traditionally employed in human creativity research (?, ?, ?). However, the task proved to be relatively easy for a large language model like GPT, which consistently outperformed humans alone both in creativity and diversity. This does not necessarily imply that AI is intrinsically more creative than humans; rather, it highlights the limitations of using tasks designed for human cognition as benchmarks for AI creativity. AI’s advantage may stem from its ability to rapidly generate highly novel yet semantically coherent text, a skill optimized by its training on vast datasets (?, ?, ?). Conversely, human creativity and creative writing is influenced by memory constraints, cognitive biases, and adherence to implicit narrative conventions (?, ?).
@@
Our study aimed to simulate social-naturalistic dynamics within experimental societies. However, our approach diverges from how ecologically valid social interactions unfold in the real world. Unlike our experiments, where participants engaged only once, individuals in real-world settings continuously interact with each other over time. Additionally, real-world exchanges are far more complex than the constrained task of creating simple stories used in our experiment. Another key limitation is the static structure of our simulated social network, which used a grid-based topology. Real social networks, such as groups of friends or work relationships, tend to have a modular structure with irregular patterns and evolving connections, allowing individuals to actively shape their interactions by forming and adjusting their social ties rather than being confined to static connections. Finally, our paradigm avoids any direct communication between participants and masks agent information, such as whether your partner is a human or an AI. Prior research suggests that awareness of interacting with an AI system may alter individual behavior, which in turn could influence collective dynamics (?, ?).
In addition, our study has several technical limitations. First, key parameters of the AI models—such as model architecture, temperature settings, and prompt design—may have influenced the results. We intend to systematically examine the impact of these variables in future work. Second, while we employed large language models (LLMs) for story generation and selection, more advanced AI agents now enable richer forms of interaction with both the environment and other agents—an aspect not addressed in the current study (?, ?). Lastly, our validation approach was relatively limited. We relied on a single measure of creativity, and participants were not provided with an explicit definition of the term. Future studies could incorporate a more comprehensive set of creativity metrics and a more robust validation framework.